The product of two proper fractions is _______ than each of the fractions that are multiplied. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The product of two proper fractions is less than each of the fractions that are multiplied. Consider the two proper fractions, 4/5 and 2/4

= (4/5) × (2/4)
= 2/5
= 0.4

Then, 0.4 is multiplied to the proper fraction less than each of the fractions that are multiplied = (4/5) × 0.4

= 0.32


Thanks